# Dissecting the SWE-Bench Leaderboards: Profiling Submitters and Architectures of LLM- and Agent-Based Repair Systems

### v1

@@ 3.4.5. Patch Generation
In the patch generation phase, approaches produce patches based on suspicious code elements identified during the localization phase (Section 3.4.3) or derived from a modification plan (Section 3.4.4). Patch generation is typically included as a distinct phase in approaches with a crafted workflow such as Agentless(xia2024agentlessdemystifyingllmbasedsoftware, ), PatchPilot(li2025patchpilot, ), and DARS (aggarwal2025darsdynamicactionresampling, ), consistent with the design of traditional APR techniques.
@@
Specialized Agents tasked with generating patches have been employed by solutions, referred to by different names, for example, Editor (Editing) Agent in Globant Code Fixer Agent, MarsCode (liu2024marscode, ), CodeR (xia2024dcoder, ), SWE-Fixer (xie2025SWEFixerTrainingopensourcellms, ), HyperAgent (phan2024hyperagent, ), IBM AI Agent SWE-1.0; Patching Agent in SpecRover (ruan2024specrovercodeintentextraction, ); and Fixer Agent in nFactorial. These agents typically receive, in addition to the suspicious code elements, the problem statement (e.g., issue description) and additional context, which may include other files, relevant code snippets, failing test cases, or —such as in AutoCodeRover (Zhang2024AutoCoderRover, )— a history of context retrieval, including invoked APIs, their results, and prior code analysis performed by other agents. In some submission, the patching agents follow a retry-loop strategy for patch generation. For instance, in AutoCodeRover(Zhang2024AutoCoderRover, ) and Alibaba LingmaAgent(ma2025alibabalingmaagent, ), if a generated patch fails to conform to the required format (e.g., an invalid diff) or cannot be syntactically applied to the original program, the agent is prompted to retry. Specialized tools for editing files and generating patches have been specifically developed in submissions such as Aider and TOOLS + Claude by Anthropic. For instance, IBM Research Agent-101 includes tools designed to prevent common errors such as incorrect indentation, missing imports, and imprecise line numbers.
@@
The patch generation phase may produce one or multiple candidate patches, which are subsequently evaluated and ranked in later phases. For instance, Agentless (xia2024agentlessdemystifyingllmbasedsoftware, ) and OpenHands (wang2024openhandsopenplatformai, ) sample multiple candidate patches per issue to maximize the chance of generating a correct fix. W&B Programmer O1 employs a parallelization strategy by generating five candidate patches simultaneously for each issue. Other submissions such as CodeFuse-AAIS and one particular from OpenHands (OpenHands + 4x Scaled (2024-02-03)) combine multiple large language models to expand the sampling space. Aider employs a cascade strategy, beginning with one LLM and, if no plausible solution is found, proceeding with another. IBM AI Agent SWE-1.0 uses multiple Editing agents —each powered by a different LLM— and a Judge agent responsible for analyzing and selecting among the resulting patches.
@@
Multiple patches can result from varying the configuration, the application of different models and strategies. For example, SuperCoder produces different solutions by sampling with different temperature settings. Similarly, SIMA submission deploys multiple agents, each configured with a distinct temperature value. Moreover, multiple patches can also result from the creation and evaluation of different execution trajectories. For example, DARS (aggarwal2025darsdynamicactionresampling, ) generates multiple patches through its Expansion mechanism, which branches the trajectory at key decision points based on prior trajectory history and execution feedback from previous attempts. Nevertheless, the generation of multiple patches can be fostered in the earlier stages of the repair pipeline. For example, PatchPilot(li2025patchpilot, ) encourages patch diversity at the planning stage (Section 3.4.4) by designing three types of prompts for plan generation. These prompts explicitly guide the LLM to produce patching plans with different focuses, for instance, generating a minimal patch with the smallest possible modifications.
@@ 3.4.6. Patch Verification
Syntax checkers or linters —static analysis tools that detect stylistic issues, potential bugs, and suspicious code constructs— are used in several submissions, including Alibaba Lingma Agent, Amazon Q Developer Agent, and Nemotron-Cortexa. Different types of tools are used across submissions to assist in patch validation. For example, Aegis and PatchPilot execute Flake8222222https://flake8.pycqa.org/en/latest/, a command-line tool that checks Python code against coding style (PEP 8), detects programming errors, and flags overly complex constructs. Artemis Agent invokes external utilities such as profilers (e.g., VTune [^23] https://www.intel.com/content/www/us/en/developer/tools/oneapi/vtune-profiler.html, Speedscope [^24] https://github.com/jlfwong/speedscope) and security analyzers (e.g., SonarQube [^25] https://github.com/SonarSource/sonarqube). CodeStory Midwit Agent relies on Pyright [^26] https://github.com/microsoft/pyright to gather diagnostics such as missing imports. Factory Code Droid extends validation further by employing not only linters and static analyzers, but also debuggers. SWE-agent integrates a linter into the edit function to automatically alert the agent of any mistakes introduced during file editing (yang2024sweagentagent, )
@@
Syntax or linter errors can be fed back into the generation phase to produce revised code that avoids the detected issues, as implemented in submissions such as SWE-agent, AppMap Navie, and CodeStory Midwit Agent. Notably, Aider prompts the LLM with linting errors represented in abstract-syntax-tree format, allowing it to display the relevant code context for each error. This enables the model to better understand the problem and make the appropriate corrections.
@@
Dynamic verification is also used to assess the plausibility of candidate patches. Traditional APRs, such as GenProg (LeGoues2011genprog, ) pioneered on the usage of test cases for semantically candidate validating patches. The majority of the submissions to SWE-Bench execute regression test and, when available, also run the issue reproduction test or script (Section 3.4.2). Submissions vary in their test execution strategies during patch validation. While some, like AppMap Navie, execute all the test cases from the repository, others, such as AgentScope, limit execution to a selected subset. Aegis, for example, filters tests based on filename similarity between test files and the modified files. In contrast, Agentless begins with a full test run and subsequently narrows the test set using LLM-guided selection. Additionally, discussed in Section 3.4.2, certain approaches (e.g., CodeShellTester) enhance patch validation by synthesizing new test cases.
@@
Patch refinement is an additional phase included in some approaches following validation. For example, PatchPilot (li2025patchpilot, ) explicitly incorporates a Refinement stage in its workflow, where the current patch is iteratively improved based on validation feedback. SpecRover (ruan2024specrovercodeintentextraction, ) introduces a Review agent that generates Reviewer Feedback —a summary derived from inferred specifications explaining the rationale behind the patch correctness judgment. This feedback is then passed to the Patching agent to guide the generation of an improved patch.
3.4.7. Ranking (Patch selection)
@@
The last phase in usually the selection of a single patch to be presented to the developer, and in the case of experimentation on SWE-Bench, to be submitted for further evaluation (e.g., using the bug-revealing tests not available during repair). For selecting the best patches, some approaches apply patch ranking.
@@
Majority voting is adopted by several submissions as a strategy for selecting the final patch. For example, Agentless (xia2024agentlessdemystifyingllmbasedsoftware, ) normalizes candidate patches to eliminate superficial syntactic differences, then selects the one with the highest frequency. Cortexa employs a similar voting mechanism. Bytedance MarsCode Agent normalizes patches at the AST level and leverages an LLM to vote across all candidates, selecting the patch receiving the most votes. Augment Agent also employs a simple majority voting strategy using the OpenAI o1 model: it presents the model with a list of candidate diffs and the problem statement, prompting it to pick the majority vote solution.
@@
Measuring similarity among candidate patches guides patch selection in certain submission For example, SIMA applies a vector similarity voting method that selects the patch most semantically aligned with the others based on embedding distance. CodeShellAgent clusters vectorized candidate patches using an embedding LLM and k-means, selecting the patch closest to the centroid of the largest cluster.
Prompting LLMs is also a strategy for selecting a single patch. For example, W&B Programmer O1, which parallelizes patch generation, uses OpenAI o1 as a tie-breaker to select the best patch. MASAI (arora2024masaimodulararchitecturesSE, ) employs an LLM to rank candidate patches based on the issue description and, when available, the results of repository test executions.
@@
Scoring the quality of patches is also employed by approaches such as Aider and AppMap Navie to support patch selection. For instance, Aider computes a score based on factors such as the number of edits, linting and testing error. DARS (aggarwal2025darsdynamicactionresampling, ) uses a fine-tuned Deepseek R1 as a Reviewer LLM to assign patch scores based on three predefined rubrics. These include criteria such as correctness of the fix and regression risk, each rated on a three-point scale.
@@
A combination of scoring and voting is also employed. For instance, SWE-RL, through its scaffold Agentless-Mini, first clusters patches into concept groups, where each group consists of patches that pass the same set of reproduction tests —a strategy shown to effectively reduce the candidate space (martinez2024xTestCluster, ). Next, the approach assigns a score to each group based on the number of passing tests and the number of patches it contains. Finally, it selects the consensus group with the highest score and applies majority voting within the group to determine the best patch.
@@
Training or fine-tuning models to evaluate candidate patches has also been explored. AgentScope employs a reward model during its voting stage, which takes as input a candidate patch and the summarized repair trajectory, assigning a score used to select the best patch. This reward model is obtained by fine-tuning an open-source model using data from benchmarks such as SWE-Gym(pan2024SWEGym, ) and SWE-bench-extra [^27] https://huggingface.co/datasets/nebius/SWE-bench-extra, a preliminary benchmark for SWE-re-bench(badertdinov2025swerebenchautomatedpipelinetask, ). OpenHands trains a dedicated critic model to select the best solution (wang2025OInferenceTimeScalingAndCritic, ), using agent trajectories from SWE-Gym (pan2024SWEGym, ). The model applies temporal difference (TD) learning to propagate success signals and incorporates a regression head to estimate reward values. Nebius employs a critic-guided search strategy (zainullina2025guidedsearchstrategies, ) that supports two forms of supervision: a) process supervision, which assesses the quality of individual actions and intermediate states, and b) outcome supervision, which evaluates the correctness of the complete trajectory leading to a patch. The critic is based on a fine-tuned LLaMA 3 model, trained on trajectories from SWE-agent using the SWE-bench-extra dataset developed by Nebius [^28] https://huggingface.co/datasets/nebius/SWE-agent-trajectories.
@@
Agents specialized in patch selection and ranking are also featured in several multi-agent submissions. For instance, SpecRover (casper2025aiagentindex, ) employs a Selection Agent that selects a patch based on the issue description and provides a justification for its choice. IBM AI Agent SWE-1.0 integrates a Judge Agent responsible for selecting the final patch from those proposed by multiple Editing Agents, each potentially powered by a different foundation LLM. Emergent E1 includes a Patch Adjudicator, an LLM-based agent that evaluates candidate patches and their associated trajectories to select the most suitable one based on the problem statement. Other submissions incorporating similar agents include Globant Code Fixer Agent (Critic Agent), nFactorial (Scorer Agent), and Bracket.sh (Judge Agent).

### v2

@@ 3.3.5. Patch Generation
In the patch generation phase, approaches produce patches based on suspicious code elements identified during the localization phase (Section 3.3.3) or derived from a modification plan (Section 3.3.4). Patch generation is typically included as a distinct phase in approaches with a crafted workflow such as Agentless (xia2024agentlessdemystifyingllmbasedsoftware, ) (40.67% on Lite, 50.8% on Verified), PatchPilot (li2025patchpilot, ) (41.33% on Lite, 64.6% on Verified), and DARS (aggarwal2025darsdynamicactionresampling, ) (47% on Lite), consistent with the design of traditional APR techniques. In this phase, submissions either rely on foundation models —as in the mentioned approaches— or use fine-tuned models. This is the case for Co-PatcherR (tang2025copatchercollaborativesoftwarepatching, ) (46% on Verified), which enhances generation by combining two LLM-driven sub-tasks: patch generation, where the model receives the issue description and identified root causes to produce a candidate patch, and patch critique, where the model evaluates a given patch and suggests corrections if needed. Using Qwen-2.5-Coder-14B as the base model, the generation LLM is finetuned on reasoning-chain distillation data produced by Claude-3.7-Sonnet. Notably, in Co-PatcherR, a single model is trained for both localization (See 3.3.3) and generation tasks, as both require similar capabilities—interpreting the issue and understanding the codebase.
@@
A number of solutions use specialized agents tasked with generating patches. These agents are referred to using different names; for example, Editor (Editing) Agent in Globant Code Fixer Agent, MarsCode (liu2024marscode, ), CodeR (xia2024dcoder, ), SWE-Fixer (xie2025SWEFixerTrainingopensourcellms, ), HyperAgent (phan2024hyperagent, ), IBM AI Agent SWE-1.0; Patching Agent in SpecRover (ruan2024specrovercodeintentextraction, ); and Fixer Agent in nFactorial. These agents typically receive, in addition to suspicious code elements, the problem statement (e.g., issue description) and additional context, which may include other files, relevant code snippets, failing test cases, or —such as in AutoCodeRover (Zhang2024AutoCoderRover, )— a history of context retrieval, including invoked APIs, their results, and prior code analysis performed by other agents. In some submission, patching agents follow a retry-loop strategy for patch generation. For instance, in AutoCodeRover(Zhang2024AutoCoderRover, ) and Alibaba LingmaAgent(ma2025alibabalingmaagent, ), if a generated patch fails to conform to the required format (e.g., an invalid diff) or cannot be syntactically applied to the original program, the agent is prompted to retry.
Authors from some submissions, such as Aider (26.33% on Lite) and TOOLS + Claude (73.2% on Verified), designed and developed specialized tools for editing files and generating patches, or as the case of IBM Research Agent-101 to prevent common errors such as incorrect indentation, missing imports, and imprecise line numbers.
Similarly to some traditional APR approaches such as (Bach2016History, ), which leverage historical repair data, ExpeRepair (mu2025experepairdualmemoryenhancedllmbased, ) (60.33% on Lite) adopts a multi-agent architecture enhanced with a dual-memory mechanism: 1) episodic memory, which stores concrete and reusable demonstrations of past repairs, and 2) semantic memory, which captures high-level, generalizable insights (e.g., repair strategies). Two of ExpeRepair’s agents, the Patch Agent and the Test Agent, leverage this memory system when generating patches and tests, respectively. They retrieve relevant demonstrations (patches or tests) from the episodic memory, and use summarized natural language insights from the semantic memory to guide their reasoning with general repair strategies.
@@
The patch generation phase may produce one or multiple candidate patches, which are subsequently evaluated and ranked in later phases. For instance, Agentless (xia2024agentlessdemystifyingllmbasedsoftware, ) and OpenHands (wang2024openhandsopenplatformai, ) sample multiple candidate patches per issue to maximize the chance of generating a correct fix. W&B Programmer O1 (64.6% on Verified) employs a parallelization strategy by generating five candidate patches simultaneously for each issue. Other submissions such as TRAE (75.2% on Verified), CodeFuse-AAIS (35.67% on Lite) and one particular from OpenHands -4x Scaled (2024-02-03)- (60.8% on Verified) combine multiple large language models to expand the sampling space. Aider (26.33% on Lite) employs a cascade strategy, beginning with one LLM and, if no plausible solution is found, proceeding with another. IBM AI Agent SWE-1.0 (23.67% on Lite) uses multiple Editing agents —each powered by a different LLM— and a Judge agent responsible for analyzing and selecting among the resulting patches. Devlo (70.2% on Verified) employs multiple LLMs with “diverse training data, inductive biases, and reasoning habits”. According to the authors, this diversity allows Devlo to surface what one model might miss, triangulate more accurate solutions, and gain confidence when models independently propose similar fixes.
Multiple patches can result from varying the configuration, the application of different models and strategies. For example, SuperCoder (34% on Lite) produces different solutions by sampling with different temperature settings. In ExpeRepair (60.33% on Lite) , the Patch Agent performs multiple samplings per iteration, with the temperature increasing at each step to promote greater diversity among candidate patches. KGCompass (yang2025enhancing, ) (46% on Lite) employs a hybrid sampling strategy, combining deterministic (temp = 0) and exploratory (temp ¿ 0) sampling methods.
@@
Similarly, SIMA (27.67% on Lite) submission deploys multiple agents, each configured with a distinct temperature value. Moreover, multiple patches can also result from the creation and evaluation of different execution trajectories. For example, DARS (aggarwal2025darsdynamicactionresampling, ) (47% on Lite) generates multiple patches through its Expansion mechanism, which branches the trajectory at key decision points based on prior trajectory history and execution feedback from previous attempts. Nevertheless, the generation of multiple patches can be fostered in the earlier stages of the repair pipeline. For example, PatchPilot (li2025patchpilot, ) (41.33% on Lite, 64.6% on Verified) encourages patch diversity at the planning stage (Section 3.3.4) by designing three types of prompts for plan generation. These prompts explicitly guide the LLM to produce patching plans with different focuses, for instance, generating a minimal patch with the smallest possible modifications. Codev (49% on Lite) launches multiple rollouts based on the number of files identified as ‘interesting’ by the fault localization stage (see Section 3.3.3). For each rollout, it generates a distinct development plan to introduce variance and explore alternative solutions. Zencoder (70% on Verified) tackles each SWE-bench task by running four agents in parallel, each equipped with a distinct LLM (Claude Sonnet 3.7 or o4-mini) and a unique set of diagnostic, search, editing, and reasoning tools.
@@ 3.3.6. Patch Verification
Syntax checkers or linters —static analysis tools that detect stylistic issues, potential bugs, and suspicious code constructs— are used in several submissions, including Alibaba Lingma Agent (33% on Lite, 28.8% on Verified), Amazon Q Developer Agent (29.67% on Lite, 65.4% on Verified), and Nemotron-CORTEXA (68.2% on Verified). Different types of tools are used across submissions to assist in patch validation. For example, Aegis (30.33% on Lite) and PatchPilot (41.33% on Lite, 64.6% on Verified) execute Flake8202020https://flake8.pycqa.org/en/latest/, a command-line tool that checks Python code against coding style (PEP 8), detects programming errors, and flags overly complex constructs. Artemis Agent (32% on Verified) invokes external utilities such as profilers (e.g., VTune [^21] https://www.intel.com/content/www/us/en/developer/tools/oneapi/vtune-profiler.html, Speedscope [^22] https://github.com/jlfwong/speedscope) and security analyzers (e.g., SonarQube [^23] https://github.com/SonarSource/sonarqube). CodeStory Midwit Agent (62.2% on Verified) relies on Pyright [^24] https://github.com/microsoft/pyright to gather diagnostics such as missing imports. Factory Code Droid ( 31.33% on Lite, 37% on Verified) extends validation further by employing not only linters and static analyzers, but also debuggers. SWE-Agent (56.67% on Lite, 66.6% on Verified) integrates a linter into the edit function to automatically alert the agent of any mistakes introduced during file editing (yang2024sweagentagent, )
@@
Syntax or linter errors can be fed back into the generation phase to produce revised code that avoids the detected issues, as implemented in submissions such as SWE-agent, AppMap Navie, and CodeStory Midwit Agent. Notably, Aider (26.33% on Lite) prompts the LLM with linting errors represented in abstract-syntax-tree format, allowing it to display the relevant code context for each error. This enables the model to better understand the problem and make the appropriate corrections.
@@
Dynamic verification is also used to assess the plausibility of candidate patches. Traditional APRs, such as GenProg (LeGoues2011genprog, ) pioneered on the usage of test cases for semantically candidate validating patches. The majority of the submissions to SWE-Bench execute regression test and, when available, also run the issue reproduction test or script (Section 3.3.2). Submissions vary in their test execution strategies during patch validation. While some, like AppMap Navie or Devlo, execute all the test cases from the repository, others, such as AgentScope (63.4% on Verified), limit execution to a selected subset. Aegis (30.33% on Lite), for example, filters tests based on filename similarity between test files and the modified files. In contrast, Agentless (40.67% on Lite, 50.8% on Verified) begins with a full test run and subsequently narrows the test set using LLM-guided selection. Additionally, discussed in Section 3.3.2, certain approaches (e.g., CodeShellTester (31.33% on Lite)) enhance patch validation by synthesizing new test cases.
@@
Patch refinement is an additional phase included in some approaches following validation. For example, PatchPilot (li2025patchpilot, ) (41.33% on Lite, 64.6% on Verified) explicitly incorporates a Refinement stage in its workflow, where the current patch is iteratively improved based on validation feedback. SpecRover (ruan2024specrovercodeintentextraction, ) (51.6% on Verified) introduces a Review agent that generates Reviewer Feedback —a summary derived from inferred specifications explaining the rationale behind the patch correctness judgment. This feedback is then passed to the Patching agent to guide the generation of an improved patch. In ExpeRepair (60.33% on Lite), patch refinement is performed during the Validation phase. In this phase, the Patch Agent improves the solution by addressing edge case handling, regression risks, and adherence to language-specific best practices, while the Test Agent enhances the test suite with additional cases targeting boundary conditions and corner cases.
LLM-as-judge have be also employed by submissions, eventually after other validations. For instance, CoDev (49% on Lite), once a solution passes the test script, it undergoes regression testing. If a regression test fails, a reasoning model assesses whether the failure is still relevant or if the test should be updated in order to ensure that valid solutions are not wrongly discarded due to outdated tests.
3.3.7. Patch Ranking and Selection
@@
The last phase is usually the selection of a single patch to be presented to the developer, and in the case of experimentation on SWE-Bench, to be submitted for further evaluation (e.g., using the bug-revealing tests not available during repair). For selecting the best patches, some approaches apply patch ranking.
@@
Majority voting is adopted by several submissions as a strategy for selecting the final patch. For example, Agentless (xia2024agentlessdemystifyingllmbasedsoftware, ) (40.67% on Lite, 50.8% on Verified) normalizes candidate patches to eliminate superficial syntactic differences, then selects the one with the highest frequency. A similar majority-based strategy is adopted by Nemotron-CORTEXA (sohrabizadeh2025nemotroncortexa, ) (68.2% on Verified). Bytedance MarsCode Agent (39.33% on Lite, 50% on Verified) normalizes patches at the AST level and leverages an LLM to vote across all candidates, selecting the patch receiving the most votes. Augment Agent (70.4% on Verified) also employs a simple majority voting strategy using the OpenAI o1 model: it presents the model with a list of candidate diffs and the problem statement, prompting it to pick the majority vote solution. TRAE (75.2% on Verified) initially adopted the majority voting from Augment Agent, but discontinued its use after observing a decline in performance as the sampling space increased. Instead, TRAE is equipped with a Selector agent with two stages: (1) Syntax-Based Voting: the agent first performs syntax-based voting by clustering patches based on AST equivalence and selecting the most frequent variant. Then, the agent verifies whether the syntax-voted patch behaves as expected using the issue description and a set of tools (e.g., bash, file editor) (2) Multi-Selection Voting: If the agent in the previous phase cannot determine the correctness of the selected patch, the process proceeds by first deduplicating the candidate patches. Then, one or more Selector agents are tasked with selecting the most likely correct patch, and the final output is determined by aggregating their votes.
@@
Measuring similarity among candidate patches guides patch selection in certain submission For example, SIMA (27.67% on Lite) applies a vector similarity voting method that selects the patch most semantically aligned with the others based on embedding distance. CodeShellAgent (44.2% on Verified) clusters vectorized candidate patches using an embedding LLM and k-means, selecting the patch closest to the centroid of the largest cluster.
Prompting LLMs is also a strategy for selecting a single patch. For example, W&B Programmer O1 (64.6% on Verified), which parallelizes patch generation, uses OpenAI o1 as a tie-breaker to select the best patch. Zencoder (70% on Verified) employs o3-mini to select the best patch from the candidates generated by four agents running in parallel (See 3.3.5). MASAI (arora2024masaimodulararchitecturesSE, ) (27.33% on Lite, 32.6% on Verified) employs an LLM to rank candidate patches based on the issue description and, when available, the results of repository test executions. In the initial submission (2025-04-25) Refact.ai Agent included a specialized tool, deep_analysis, which can be invoked during the repair workflow. This tool follows a three-step process: after generating an initial patch, it performs a Critique step that identifies weaknesses, limitations, or bugs in the generated code. Finally, in the Refinement step, the patch is improved based on the feedback from the Critique phase. This tool is backed by a reasoning model o4-mini that handles the cognitive load of problem. However, in a later submission, the authors removed the Critique step, and the tool now only generates a candidate patch. They found that the approach performs better when it simply runs tests and determines the next steps based on the results. ExpeRepair (60.33% on Lite) has a dedicated review agent, which selects the final patch based on criteria such as correctness, code style, and adherence to best practices.
@@
Scoring the quality of patches is also employed by approaches such as Aider and AppMap Navie to support patch selection. For instance, Aider (26.33% on Lite) computes a score based on factors such as the number of edits, linting and testing error. DARS (aggarwal2025darsdynamicactionresampling, ) (47% on Lite) uses a fine-tuned Deepseek R1 as a Reviewer LLM to assign patch scores based on three predefined rubrics. These include criteria such as correctness of the fix and regression risk, each rated on a three-point scale.
@@
A combination of scoring and voting is also employed. For instance, SWE-RL (41.2% on Verified), through its scaffold Agentless-Mini, first clusters patches into concept groups, where each group consists of patches that pass the same set of reproduction tests —a strategy shown to effectively reduce the candidate space (martinez2024xTestCluster, ). Next, the approach assigns a score to each group based on the number of passing tests and the number of patches it contains. Finally, it selects the consensus group with the highest score and applies majority voting within the group to determine the best patch. Similarly, Codev (49% on Lite) applies a cross-validation algorithm, inspired the technique xTestCluster (martinez2024TestCluster, ): given patches and test scripts generated by different trajectories, the technique validates a patch from one trajectory against the generated test from the other trajectories. Codev then assigns scores based primarily on the results of the generated test scripts and regression tests.
@@
Training or fine-tuning models to evaluate candidate patches has also been explored. AgentScope (63.4% on Verified) employs a reward model during its voting stage, which takes as input a candidate patch and the summarized repair trajectory, assigning a score used to select the best patch. This reward model is obtained by fine-tuning an open-source model using data from benchmarks such as SWE-Gym(pan2024SWEGym, ) and SWE-bench-extra [^25] https://huggingface.co/datasets/nebius/SWE-bench-extra, a preliminary benchmark for SWE-re-bench(badertdinov2025swerebenchautomatedpipelinetask, ). OpenHands trains a dedicated critic model to select the best solution (wang2025OInferenceTimeScalingAndCritic, ), using agent trajectories from SWE-Gym (pan2024SWEGym, ). The model applies temporal difference (TD) learning to propagate success signals and incorporates a regression head to estimate reward values. Nebius (40.6% on Verified) employs a critic-guided search strategy (zainullina2025guidedsearchstrategies, ) that supports two forms of supervision: a) process supervision, which assesses the quality of individual actions and intermediate states, and b) outcome supervision, which evaluates the correctness of the complete trajectory leading to a patch. The critic is based on a fine-tuned LLaMA 3 model, trained on trajectories from SWE-agent using the SWE-bench-extra dataset developed by Nebius [^26] https://huggingface.co/datasets/nebius/SWE-agent-trajectories.
@@
Agents specialized in patch selection and ranking are also featured in several multi-agent submissions. For instance, SpecRover (casper2025aiagentindex, ) (51.6% on Verified) employs a Selection Agent that selects a patch based on the issue description and provides a justification for its choice. IBM AI Agent SWE-1.0 (23.67% on Lite) integrates a Judge Agent responsible for selecting the final patch from those proposed by multiple Editing Agents, each potentially powered by a different foundation LLM. Emergent E1 (57.2% on Verified) includes a Patch Adjudicator, an LLM-based agent that evaluates candidate patches and their associated trajectories to select the most suitable one based on the problem statement. Other submissions incorporating similar agents include Globant Code Fixer Agent (Critic Agent), nFactorial (Scorer Agent), and Bracket.sh (Judge Agent).